Etna (Sicily, Italy): Activity continues with little changes, at fluctuating and overall slowly decreasing intensity, reflected also by the tremor signal. Last evening, liquid lava spattering from the central vent of the New SE crater occurred at rapid succession reaching heights of 50-150 m and often throwing incandescent material to the upper outer flanks of the cone.
Lava flows continued to descend through the channel on the eastern side of the cone and spread into several channels on the upper half of the Valle del Bove wall, but not extending beyond the fronts emplaced during yesterday night.
The situation this morning is similar, but the spattering at the summit has decreased.
More at Volcano Discovery: Volcanoes Today, 17 Jun 2014: Etna